OpenStudy (anonymous):

What is 5/9 minus 3/4? I got -7/36. Is this wrong?

OpenStudy (anonymous):

You are absolutely right...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[\huge \color{blue} {\frac{5}{9} - \frac{3}{4} = \frac{20-27}{36} = \frac{-7}{36}}\]

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

Copy this: frac{}{} Now paste them in equation editor and fill the value inside the brackets...